OpenStudy (anonymous):

In a proportion, where a ≠ 0 and b ≠ 0, which of the following are equivalent to a/2 = 3/b ? Help! I will medal and fan. :)

OpenStudy (anonymous):

@ganeshie8 @Taylor<3sRin

OpenStudy (taylor<3srin):

Eh, Math isn't my strongest subject.

OpenStudy (anonymous):

that's ok, just thought I'd try :)

OpenStudy (anonymous):

@ganeshie8 @texaschic101

OpenStudy (texaschic101):

proportions are nothing but equivalent fractions. a/2 = 3/b 6/2 = 3/1 3 = 3 (correct)

OpenStudy (anonymous):

what in the nine hells...never mind, the question disappeared for awhile, couldn't find it Anyway, thnx for your help. I get it now.

OpenStudy (texaschic101):

no problem :)